The solubility product constant for SrCO 3 is 3.5 * 10 ^ - 8 at 25°C. The solubility of SrCO 3 is ______ M
Answer #1

SrCO3(s) = Sr+2(aq) + CO3-2(aq)

Ksp = [Sr+2][CO32-]

Note = solid and liquid state does not count in equilibrium calculations

3.5*10-8 = X2

squareroot(3.5*10-8) = X

hence X = 1.9*10-4M

hence solubility of SrCO3 is 1.9*10-4M